OpenStudy (anonymous):

Limit Question: Okay, so using the DEFINITION of a limit prove x^2-2x-1=-2. Here's the issue I have with it. I understand the formula completely, I get how to do it, but in my advanced math class I have to create a graph of limits using Maple and I can't figure out how to do it when you break it up and you get |x-1||x-1|...it just doesn't make sense to me. Any help? Also, if you don't understand I'll try to make it clearer lol

OpenStudy (agent0smith):

I'm a bit confused here... x^2-2x-1=-2 is only valid for x = 1. But x^2-2x-1 can also be equal to any value greater/equal to -2. How do you "prove" that x^2-2x-1=-2 when it's not even true unless x=1? Maybe I need to look at the definition of a limit again...

OpenStudy (anonymous):

Here lemme explain it a little bit better: lim(x->1): x^2 - 2x - 1 = -2 So, when you prove it it turns into this: |(x^2 - 2x - 1) + 2|=|x^2 - 2x + 1|=|x-1||x-1| What I'm confused about it the next step in terms of a formal definition. Usually when you have a quadratic equation you have (-) and a (+) number so you have to bound the higher value but the solutions in this case are the same! So I have no idea how to continue. Does this make more sense?

OpenStudy (anonymous):

Normally, I'd try to solve it with two cases. One case being when some \[\epsilon\] is less than my biggest bound (I use 1/2) and then another case being when some epsilon is greater than that bound. But when it's the same number, it's confusing.

OpenStudy (anonymous):

i will help u in a bit, i just have a phone call, ur problem is sorted

OpenStudy (anonymous):

awesome! thanks!

OpenStudy (jtvatsim):

doesn't |x-1||x-1| < e just become |x-1|^2 < e, so |x-1| < sqrt(e), then 1 - sqrt(e) < x < 1 + sqrt(e) ?

OpenStudy (anonymous):

In all of my examples I would never sqrt(e)....I'm very confused with why you did that. I don't even understand how you'd continue, BUT I will try it first to see if I get anything helpful before I criticize :)

OpenStudy (jtvatsim):

yeah, I know it seems weird, but since the error term is always greater than 0 taking the square root is mathematically sound at least.

OpenStudy (anonymous):

I tried it but I can't see how it works with the formula. |x-1|^2 < delta |x-1| < sqrt(delta) So |x-1|*sqrt(delta) Idk that just doesn't seem right at all...

OpenStudy (jtvatsim):

Well, I don't think we would use |x-1|^2 < delta. For the proof of a limit don't we use: |x - limit| < delta? So since our limit is 1, we have |x - 1| < delta. But we can just choose delta = sqrt(e) since we know that |x-1| < sqrt(e). That proves that there exists a delta that will satisfy the error formula.

OpenStudy (jtvatsim):

That is, for any given error bound, just choose the square root of the error for delta and you will stay within the bounds.

OpenStudy (anonymous):

But we don't know |x-1| < sqrt(e) yet....because you can't just say that delta = sqrt(e)...can you? Okay, this squaring this is confusing me so much now lol

OpenStudy (jtvatsim):

OK, we know that |x - 1| < sqrt(e) because we prove that as follows: By definition, |Function - Limit Value| < e |x^2 -2x - 1+ 2| < e |x^2 - 2x + 1| < e |(x-1)(x-1)|<e |(x-1)^2|<e |x-1|^2<e |x-1|<sqrt(e) This proves that for our limit to work it must be such that |x-1| < sqrt(e).

OpenStudy (jtvatsim):

Now, we can ensure that this is the case by turning to the second part of the definition: |x - Limit| < delta |x - 1| < delta We need only find ANY delta that will make the equation work. So we know that if |x-1| < sqrt(e) it will work. So just pick a delta that works, namely, delta = sqrt(e). To prove that a limit exists, you only need to show that there exists SOME, ANY value of delta that satisfies the error equation.

OpenStudy (anonymous):

Okay, well I'm exhausted. But I will look at this again tomorrow. Thanks for your help!

OpenStudy (jtvatsim):

no problem, good luck!

OpenStudy (anonymous):

x^2-2x-1 =X^2 -2x +1 -2 =(x-1)^2 -2 here (x-1)^2 is always positive hence the least value of (x-1)^2 is 0 hence the least value of the expression x^2-2x-1 is -2

OpenStudy (anonymous):

The work that @jtvatsim has shown is absolutely correct. However, I must point out that the definition of the limit does not operate in that way. Your proof would be sufficient for ordinary professors, but possibly not all. The reason is that by definition: lim (x =>x1) f(x) = L if and only if for every epsilon > 0, there exists a delta > 0 such that |f(x) - L| < epsilon whenever |x-x1|<delta. The keyword here is whenever. A whenever B is not A implies B but rather B implies A. Your work is merely a way of estimating delta in terms of epsilon. the complete, foolproof proof requires that you show |x-x1| < delta => |f(x)-L| < epsilon when for delta as some function of epsilon.

OpenStudy (anonymous):

you want an \(\epsilon, \delta\) proof right?

OpenStudy (anonymous):

and you have reached the step where you want a \(\delta\) so that \[|x^2-2x+1|<\epsilon\]

OpenStudy (anonymous):

and have seen that this is the same as wanting \[|x-1||x-1|=|x-1|^2<\epsilon\] now you pick the \(\delta\) in terms of \(\epsilon\) and work the algebra backwards is all you need to do. that is, pick \(\delta=\sqrt{\epsilon}\)

OpenStudy (anonymous):

too much work if \(\delta=\sqrt{\epsilon}\) then given \(|x-1|<\sqrt{\epsilon}\) you have \[|x-1||x-1|<\sqrt{\epsilon}^2=\epsilon\] whic is all you need

OpenStudy (anonymous):

all the work is finding the appropriate \(\delta\) in terms of \(\epsilon\) the "proof" is just working the algebra backwards

OpenStudy (anonymous):

This is my "finished" proof. I think I finally got it. 'Think' being the key word ;)

OpenStudy (anonymous):

\[Choose \delta = \sqrt{\epsilon}\] By the definition of the limit, \[|x-1| < \delta = \sqrt{\epsilon}\] which implies \[|x-1|^{2} < \epsilon\] which implies \[|(x-1)^{2}| = |x ^{2}-2x+1| = |(x ^{2}-2x -1) - (-2)| = |f(x) - L|< \epsilon\]. QED. There is no need to consider cases.

OpenStudy (anonymous):

The only reason I'm doing cases is because that's what my professor told me to do. I don't understand why I'd 'have' to but I'm doing it to appease her.

OpenStudy (anonymous):

Do you know why she is doing so? That's actually pretty odd. There is no reason why 1/2 should be important. epsilon = 1/4 has virtually no significance in this limit proof.

OpenStudy (jamesj):

In any case, your case 2 proof isn't quite right as you need to replace one of the equality signs with inequality. The formal proof of this limit doesn't require cases. Many problems do, but this isn't one of them.